# Neutrino Oscillation Prospects with a Dual-Baseline Beam from BNL to SNOLAB and SURF

### v1

@@ Abstract
The Electron–Ion Collider (EIC) is a next-generation accelerator primarily designed to study the internal structure of nucleons through high-precision electron–hadron collisions. In this work, we explore the feasibility of employing a 1 MW fraction of the EIC proton beam to generate a high-intensity GeV-scale neutrino beam for long-baseline oscillation studies. We have simulated proton–target interactions and optimize the resulting neutrino fluxes for water-based liquid scintillator (WbLS) detectors located at distinct baselines of 900 km and at 2900 km. Oscillation analyses performed with GLoBES show that extended baselines allow access to multiple oscillation maxima, significantly enhancing sensitivity to leptonic CP violation. The study also examines the interplay between matter effects and the intrinsic CP violating phase in shaping observable asymmetries. These results suggest that the EIC proton beam could provide a novel and complementary source for precision neutrino physics, extending the scientific reach of the EIC program.
@@ 1 Introduction
Neutrinos continue to occupy a central place in the search for physics beyond the Standard Model. The discovery that they oscillate between flavors, implying non-zero mass and lepton mixing, has been one of the most important breakthroughs in particle physics in the past two decades Super-Kamiokande:1998kpq ; SNO:2001kpb ; nobel2015 . Standard 3-flavour neutrino oscillation is governed by two mass-squared differences, $\Delta m_{31}^{2}$ and $\Delta m_{21}^{2}$, the three mixing angles, $\theta_{12},\theta_{13},\theta_{23}$ and the CP phase, $\delta_{\text{CP}}$ Pontecorvo:1957qd ; Pontecorvo:1957cp ; Gribov:1968kq ; Maki:1962mu . Oscillation experiments have already determined two distinct mass-squared differences ($\Delta m_{21}^{2},|\Delta m_{31}^{2}|$) and two mixing angles ($\theta_{12},\theta_{13}$), and the global analyses of available neutrino data Capozzi:2018ubv ; 10.5281/zenodo.4726908 ; deSalas:2020pgw ; nufit_globalfit ; Esteban:2024eli is consistent with the standard $3\nu$ scenario. However, several fundamental puzzles in the neutrino oscillation sector still remain unresolved. Chief among these are the size and nature of CP violation in the lepton sector; the ordering of the neutrino mass states (normal versus inverted hierarchy); the correct octant of the atmospheric mixing angle $\theta_{23}$, as well as the possibilities of the presence of several non-standard oscillation effects SajjadAthar:2021prg ; Arguelles:2022tki . Addressing these questions is essential not only for completing our picture of neutrino physics, but also for potentially uncovering deeper mysteries such as the origin of matter-antimatter asymmetry in the universe or the process of generation of (tiny) neutrino mass etc.
@@
The upcoming Electron-Ion Collider (EIC) Accardi:2012qut ; Willeke:2021ymc at Brookhaven National Laboratory offers an opportunity to establish such a complementary program. Although the EIC’s primary mission is to map the three-dimensional structure of nucleons and nuclei through high-energy electron–proton and electron–ion collisions, its accelerator complex is also capable of delivering high-energy polarized proton beams. These beams, when directed at a fixed target, can generate a conventional pion-decay neutrino beam similar to the NuMI facility at Fermilab, but at substantially higher proton energies (up to 275 GeV). The resulting neutrino flux would be highly collimated and broadband, offering unique advantages for long-baseline oscillation studies. The neutrino beam could be directed toward two far detectors, - one at the SNOLAB site in Canada (with a baseline of 900 km), and one at the SURF site in USA (with a baseline of 2900 km)  [^1] Note that the SNOLAB site is currently set to host the SNO+ experiment SNO:2021xpa , which, in addition to be having a water phase SNO:2018ydj and a liquid scintillator phase SNO:2020fhu , is also proposed to have a hybrid water based liquid scintillator SNOLAB:2025vfr . The SURF site, on the other hand is set to host the above mentioned DUNE project. The shorter baseline allows precision measurements near the first oscillation maximum with reduced matter effects. The longer baseline allows probing more than one oscillation cycles (which is otherwise difficult in LBL experiments), as well as studying matter induced modifications, - thereby offering a powerful probe to CP violation and mass hierarchy. Together, the two baselines allow for simultaneous exploration of different oscillation regimes, capturing both the first and second oscillation maxima with complementary strengths. Comparisons between detectors at different baselines also enable cross-checks that reduce systematic uncertainties. Finally, the EIC’s ability to deliver polarized proton beams could open entirely new avenues, such as studying spin-dependent effects in meson production and their impact on neutrino fluxes which is a possibility not accessible in existing neutrino programs.
@@
In this work, we explore the feasibility and physics potential of such an EIC-based dual-baseline neutrino experiment. We begin in Sec. 2 by outlining the motivation and advantages of using the EIC proton beam for neutrino production. Sec. 3 presents flux calculations for the proposed baselines, while Sec. 4 analyzes oscillation probabilities and their dependence on matter effects and CP violation. Sec. 5 discusses expected event spectra at the SNOLAB and SURF detectors. Sec. 6 provides preliminary $\Delta\chi^{2}$ sensitivity estimates for CP violation. We conclude in Sec. 7 with a discussion of the broader implications, design considerations, and future directions for integrating such a neutrino program into the EIC infrastructure.
@@ 2 Motivation
The EIC is a state-of-the-art particle accelerator facility being built to solve long-standing mysteries in nuclear physics. Its primary mission is to precisely map the internal 3D structure of nucleons and nuclei by colliding high-energy polarized electron beams with polarized proton and ion beams. By doing so, it will explore the dynamics of quarks and gluons, the fundamental constituents of visible matter, and seek to understand the origin of the proton’s spin and mass. While designed for electron-ion collisions, the EIC’s powerful proton accelerator complex provides capabilities that can be harnessed for a world-class neutrino program. Comparison with the proton beam from the Long-Baseline Neutrino Facility (LBNF) at Fermilab, which powers the DUNE experiment, highlights the unique advantages and complementarities. The key parameters of the two proton beamlines are summarized in Table 1 DUNE:2020lwj ; Willeke:2021ymc .
@@
Parameter
Fermilab LBNF/DUNE EIC Collider Mode (Stored Beam) EIC Fixed-Target (Hypothetical)
@@
The EIC offers a significantly higher proton energy (up to 275 GeV) compared to LBNF’s maximum of 120 GeV. This higher energy would produce a more energetic and highly collimated neutrino beam, increasing the neutrino flux in the forward direction and extending the physics reach for very long baseline experiments  [^2] From Table 1, we note that the POT/second for the proton beam in EIC can be assumed to be $2.6\times 10^{11}\times 1160=3.016\times 10^{14}$. This number when multiplied by the proton energy of 275 GeV, gives 13.2 MW. In order to use only a part of the main proton beam at EIC for neutrino production we take a conservative assumption of an 1 MW beam.. While the instantaneous intensity of a single LBNF pulse is much higher, the EIC’s near-continuous stream of bunches could potentially deliver a substantial number of protons-on-target (POT) over a given run period.
@@
Crucially, the EIC will be the world’s only accelerator capable of delivering highly polarized proton beams at this energy scale. Repurposing this polarized beam for neutrino production would be an interesting step. It would allow for unique studies of spin-transfer in meson (pion and kaon) production, offering a new tool to probe and constrain the systematic uncertainties related to the neutrino production mechanism. In particular, given the protons in the EIC beamline can be both longitudinal and transverse polarized, there are multiple ways to study the meson production and decay processes through the controlled polarization. For example, the transverse polarized proton beam may result in modulation of the meson angular distribution. The meson angular distribution will directly impact the neutrino energy spectrum given fixed detector location. A transversely movable near detector can further explore the nature of the meson production and decay processes. This topic worth further studies in detail, which can form a separate good motivation for the proposed project. However, in this work, we will explore the benefit to the neutrino oscillation research from the super-long-baseline with unique L/E.
@@
A very specific advantage for the next-generation neutrino oscillation experiments is the broad-band beam. In the past experiments such as T2K and NOvA, they are largely counting experiments comparing the total numbers of $\nu_{e}$ and $\bar{\nu}_{e}$ appearance. With the broad-band beam, the shape information can be exclusively used to gain better sensitivity to the neutrino oscillation parameters, particularly the CP violation phase. The baseline from Brookhaven National Lab to SURF is roughly two times than the baseline from Fermilab to SURF, thus the first few $\nu_{e}$ oscillation peaks for the longer baseline are higher in terms of energy. Meanwhile, given the overall higher energy proton at the EIC beamline, we could expect more clear second and even third oscillation peaks for such a super long baseline experiment. In our calculation, we dive into the detial how the EIC baseline and energy could impact the neutrino oscillation parameter sensitivities.
@@ 3 Flux Calculation
We calculate the flux following the neutrino flux generation method in NuMI beam in Fermilab. A Geant4 model has been carried out to implement the neutrino flux simulation GEANT4:2002zbu . The 275 GeV protons from the storage ring are dumped into a pipe hitting a target. The target is assumed to be a long thin graphite with dimension of 1.5 m x 8 mm (length x radius). The resulting hadrons were focused with six horn current settings ranging from 93 kA to 593 kA with a step of 100 kA. The two horn dimensions are assumed to be 4.2 m x 0.4 m and 3.8 m x 0.48 m (length x radius) respectively.
After the horn focusing of the hadrons, a 201.7 m long pipe is simulated to allow the hardons to decay into neutrinos. The flux window at the very far-away detector cross section is taken to account for the arriving neutrinos.
@@
In Fig. 2, the simulated neutrino fluxes with 93 kA (red) and 493 kA (blue) are shown in the top left panel. The peaks match the first $\nu_{e}$ appearane peak locations. The top middle panel shows the neutrino interaction cross-section curves for various interaction channels. The top right panel shows the $\nu_{e}$ appearance oscillation probability for the two baselines. The bottom panels show the product of the flux and cross section, as well as the product of the flux, cross section and $\nu_{e}$ appearance oscillation proability. Each bottom panel corresponds to one interaction channel, namely charged-current inclusive, charged-current quasi-elastic and charged-current non-quasi-elastic from the left to right.
The CP violation phase sensitivity with our proposed detector depends on the solid curves. The final sensitivity includes the effect of absolute event rate, energy smearing and systematic uncertainties. It is obvious that for the 900 km baseline scenario, it is likely we can utilize only one appearance oscillation peak. On the contrary, for the 2900 km baseline scenario, we may use up to three oscilaltion peaks. The detector smearing may play an important to identify the oscillation peaks.
Figure 2: The top row shows the $\nu_{\mu}$ flux $\Phi_{{\nu_{\mu}}}$; $\nu_{e}$ cross-section $\frac{\sigma_{{\nu_{e}}}}{E}$; and the transition probabilities $\text{P}_{\mu e}$ at EIC-SURF baseline of 2900 km, and at EIC-SNO baseline of 900 km. The bottom row shows the relevant products of these quantities for three different types of interactions, -namely inclusive charge current ($C$), quasielastic charge current (CCQE), non-quasielastic charge current (CCNQE).

### v2

@@ Abstract
The Electron–Ion Collider (EIC) is a next-generation accelerator primarily designed to study the internal structure of nucleons through high-precision electron–hadron collisions. In this work, we explore the feasibility of employing a 1 MW fraction of the EIC proton beam to generate a high-intensity GeV-scale neutrino beam for long-baseline oscillation studies. We have simulated proton–target interactions and optimize the resulting neutrino fluxes for water-based liquid scintillator (WbLS) detectors located at distinct baselines of 900 km and at 2900 km. Oscillation analyses performed with GLoBES show that extended baselines allow access to multiple oscillation maxima, significantly enhancing sensitivity to leptonic CP violation. The study also examines the interplay between matter effects and the intrinsic CP violating phase in shaping observable asymmetries. We note that simplified systematics and no backgrounds are used in this analysis to establish the baseline physics potential. These results suggest that the EIC proton beam could provide a novel and complementary source for precision neutrino physics, extending the scientific reach of the EIC program.
@@ 1 Introduction
Neutrinos continue to occupy a central place in the search for physics beyond the Standard Model. The discovery that they oscillate between flavors, implying non-zero mass and lepton mixing, has been one of the most important breakthroughs in particle physics in the past two decades Fukuda and others (1998); Ahmad and others (2001); Kajita and McDonald . Standard 3-flavour neutrino oscillation is governed by two mass-squared differences, $\Delta m_{31}^{2}$ and $\Delta m_{21}^{2}$, the three mixing angles, $\theta_{12},\theta_{13},\theta_{23}$ and the CP phase, $\delta_{\text{CP}}$ Pontecorvo (1957a, b); Gribov and Pontecorvo (1969); Maki et al. (1962). Oscillation experiments have already determined two distinct mass-squared differences ($\Delta m_{21}^{2},|\Delta m_{31}^{2}|$) and two mixing angles ($\theta_{12},\theta_{13}$), and the global analyses of available neutrino data Capozzi et al. (2018); P. F. De Salas et al. (2021); de Salas et al. (2021); Esteban et al. ; Esteban et al. (2024) is consistent with the standard $3\nu$ scenario. However, several fundamental puzzles in the neutrino oscillation sector still remain unresolved. Chief among these are the size and nature of CP violation in the lepton sector; the ordering of the neutrino mass states (normal versus inverted hierarchy); the correct octant of the atmospheric mixing angle $\theta_{23}$, as well as the possibilities of the presence of several non-standard oscillation effects Sajjad Athar and others (2022); Argüelles and others (2023). Addressing these questions is essential not only for completing our picture of neutrino physics, but also for potentially uncovering deeper mysteries such as the origin of matter-antimatter asymmetry in the universe or the process of generation of (tiny) neutrino mass etc.
@@
The upcoming Electron-Ion Collider (EIC) Accardi and others (2016); Aschenauer et al. (2019); Abdul Khalek and others (2022a); Willeke, Ferdinand et al (2021); Abdul Khalek and others (2022b); Xu and others (2022); Adkins and others (2025) at Brookhaven National Laboratory offers an opportunity to establish such a complementary program. Although the EIC’s primary mission is to map the three-dimensional structure of nucleons and nuclei through high-energy electron-proton and electron-ion collisions, its accelerator complex is also capable of delivering high-energy polarized proton beams. These beams, when directed at a fixed target, can generate a conventional pion-decay neutrino beam similar to the NuMI facility at Fermilab, but at substantially higher proton energies (up to 275 GeV) Abdul Khalek and others (2022a); Willeke, Ferdinand et al (2021); Abdul Khalek and others (2022b). The resulting neutrino flux would be highly collimated and broadband, offering unique advantages for long-baseline oscillation studies. The neutrino beam could be directed toward two far detectors, - one at the SNOLAB site in Canada (with a baseline of 900 km), and one at the SURF site in USA (with a baseline of 2900 km)  [^1] Note that the SNOLAB site is currently set to host the SNO+ experiment Albanese and others (2021), which, in addition to be having a water phase Anderson and others (2019) and a liquid scintillator phase Anderson and others (2021), is also proposed to have a hybrid water based liquid scintillator Diamond and others (2025). The SURF site, on the other hand is set to host the above mentioned DUNE project.. The shorter baseline allows measurements of oscillation parameters near the first oscillation maximum with reduced matter effects. The longer baseline allows probing more than one oscillation cycles (which is otherwise difficult in LBL experiments), as well as studying matter induced modifications, - thereby offering a powerful probe to CP violation and mass hierarchy  [^2] The authors of Denton and Gehrlein (2025) have also studied neutrino oscillation in the context of BNL-SURF baseline, but in the context of a neutrino factory setup at BNL. The readers may also look at Denton et al. (2025) in order to understand how such a neutrino factory setup can probe non-standard neutrino interactions.. Together, the two baselines allow for simultaneous exploration of different oscillation regimes, capturing both the first and second oscillation maxima with complementary strengths. Comparisons between detectors at different baselines also enable cross-checks that reduce systematic uncertainties. Finally, the EIC’s ability to deliver polarized proton beams could open entirely new avenues, such as studying spin-dependent effects in meson production and their impact on neutrino fluxes which is a possibility not accessible in existing neutrino programs. As a first-order feasibility study, we employ simplified systematics and neglect backgrounds in this analysis.
@@
In this work, we explore the feasibility and physics potential of such an EIC-based dual-baseline neutrino experiment. We begin in Sec. 2 by outlining the motivation and advantages of using the EIC proton beam for neutrino production. Sec. 3 presents flux calculations for the proposed baselines, while Sec. 4 analyses oscillation probabilities and their dependence on matter effects and CP violation. Sec. 5 discusses expected event spectra at the SNOLAB and SURF detectors. Sec. 6 provides preliminary $\Delta\chi^{2}$ sensitivity estimates for CP violation. We conclude in Sec. 7 with a discussion of the broader implications, design considerations, and future directions for integrating such a neutrino program into the EIC infrastructure.
@@ 2 Motivation
The EIC is a state-of-the-art particle accelerator facility being built to solve long-standing mysteries in nuclear physics. Its primary mission is to precisely map the internal 3D structure of nucleons and nuclei by colliding high-energy polarized electron beams with polarized proton and ion beams. By doing so, it will explore the dynamics of quarks and gluons, the fundamental constituents of visible matter, and seek to understand the origin of the proton’s spin and mass. While designed for electron-ion collisions, the EIC’s powerful proton accelerator complex provides capabilities that can be harnessed for a world-class neutrino program. Comparison with the proton beam from the Long-Baseline Neutrino Facility (LBNF) at Fermilab, which powers the DUNE experiment, highlights the unique advantages and complementarities. The key parameters of the two proton beamlines are summarized in Table 1 Abi and others (2020a); Willeke, Ferdinand et al (2021).
@@
Parameter Fermilab LBNF/DUNE EIC Collider Mode (Stored Beam) EIC Fixed-Target (Hypothetical)
@@
The EIC offers a significantly higher proton energy (up to 275 GeV) compared to LBNF’s maximum of 120 GeV. This higher energy would produce a more energetic and highly collimated neutrino beam, increasing the neutrino flux in the forward direction and extending the physics reach for very long baseline experiments  [^3] From Table 1, we note that the POT/second for the proton beam in EIC can be assumed to be $2.6\times 10^{11}\times 1160=3.016\times 10^{14}$. This number when multiplied by the proton energy of 275 GeV, gives 13.2 MW. In order to use only a part of the main proton beam at EIC for neutrino production we take a conservative assumption of an 1 MW beam.. While the instantaneous intensity of a single LBNF pulse is much higher, the EIC’s near-continuous stream of bunches could potentially deliver a substantial number of protons-on-target (POT) over a given run period. The idea of a near-continuous stream here refers to the high-repetition pulsed structure of the proton beam at the EIC, which differs from the lower-duty-factor, widely spaced pulses used at conventional neutrino facilities. To convert such a beam into towards a fixed target or decay region for neutrino production, additional beam-handling elements, (for e.g., fast-extraction systems Kissler et al. (1979); Fraser (2018); Shibata et al. (2018), RF/microwave-based deflection or bunch-structure manipulation techniques Akre et al. (2001); Burt (2012)) - would need to be implemented. These accelerator-engineering aspects require a dedicated feasibility study and lie beyond the scope of the present paper. Here we focus solely on the baseline physics potential under idealized beam-delivery assumptions.
@@
Crucially, the EIC will be the world’s only accelerator capable of delivering highly polarized proton beams at this energy scale. Repurposing this polarized beam for neutrino production would be an interesting step. It would allow for unique studies of spin-transfer in meson (pion and kaon) production, offering a new tool to probe and constrain the systematic uncertainties related to the neutrino production mechanism. In particular, given the protons in the EIC beamline can be both longitudinal and transverse polarized, there are multiple ways to study the meson production and decay processes through the controlled polarization. For example, the transverse polarized proton beam may result in modulation of the meson angular distribution. The meson angular distribution will directly impact the neutrino energy spectrum given fixed detector location. A transversely movable near detector can further explore the nature of the meson production and decay processes.
Specifically, the transverse polarization of the proton beam induces a Single Spin Asymmetry (SSA) in the production of secondary mesons ($p^{\uparrow}C\to\pi X$), - well established in, for instance, in the data from Fermilab E704 Adams and others (1996); Bravar and others (1996), Relativistic Heavy Ion Collider (RHIC) Adams and others (2004); Arsene and others (2008); Adare and others (2014) among others. This asymmetry results in a controlled azimuthal modulation of the meson flux entering the decay volume, - giving rise to a left-right angular asymmetry. At accelerator energies the pion decay produces neutrinos that are mostly confined within a narrow cone around the pion direction, and thus this asymmetry is expected to survive in the resulting neutrinos. By measuring this angular asymmetry variation in the neutrino rate with spin-flip, one can disentangle geometric beamline systematics (which do not flip) from production physics, offering a novel in-situ constraint on the neutrino flux that is unique to the EIC. In order to get an estimate of the constraint, we first note that in our setup below, a $\mathcal{O}(10)$ kt far detector at 2900 km from EIC detects roughly $\mathcal{O}(10^{3})$ neutrino events in a year. Hence a suitably placed 11t near detector at around 1 km from the EIC source will detect roughly $N\sim\mathcal{O}(10^{5}-10^{6})$ events, - offering a statistical uncertainly of approximately $1/\sqrt{N}\sim 0.1\%$. Further, we note that the hadron production modeling is one of the principal sources of systematics in the LBL experiments Ferrari et al. (2005); Agostinelli and others (2003); Yarba (2012). Thus a model based on SSA predicting an angular asymmetry of, say, $15\%$ (see, for e.g., Adams and others (1995); Hui et al. (2004); Abelev and others (2008)) can potentially be ruled out by data showing $20.0\%\pm 0.1\%$ with a high statistical significance $\sim 50\sigma$. This would, in turn, help constrain the model parameters tightly and bring down the underlying cross-section uncertainties in the hadron production models and the resulting neutrino flux uncertainties. This topic worth further studies in detail, which can form a separate good motivation for the proposed project.
@@
However, in this work, we will explore the benefit to the neutrino oscillation research from the super-long-baseline with unique $L/E$. A very specific advantage for the next-generation neutrino oscillation experiments is the broad-band beam. In the past experiments such as T2K and NOvA, they are largely counting experiments comparing the total numbers of $\nu_{e}$ and $\bar{\nu}_{e}$ appearance. With the broad-band beam, the shape information can be exclusively used to gain better sensitivity to the neutrino oscillation parameters, particularly the CP violation phase. However, the exact shape of the beam depends on the design of the focusing horn (horn length, target position relative to the horn, as well as on horn current, among others) that determines how the mesons are transported forward and eventually decay to generate neutrinos. The upcoming DUNE with a 1300 baseline from Fermilab to SURF is set to employ the intense 1.2 MW wide band on-axis neutrino beam covering from near 1 GeV to over 10 GeV neutrino energies with a peak around 2.25-2.5 GeV. The baseline from Brookhaven National Lab to SURF is roughly two times than the baseline from Fermilab to SURF, thus the first few $\nu_{e}$ oscillation peaks for the longer baseline are higher in terms of energy. Meanwhile, given the overall higher energy proton at the EIC beamline, we could expect more clear second and even third oscillation peaks for such a super long baseline experiment. In our calculation, we dive into the detail how the EIC baseline and energy could impact the neutrino oscillation parameter sensitivities.
@@ 3 Flux Calculation
We calculate the flux following the neutrino flux generation method in NuMI beam in Fermilab. A Geant4 model has been carried out to implement the neutrino flux simulation Agostinelli and others (2003) following the LBNF beamline design report Papadimitriou and others (2016). The 275 GeV protons from the storage ring are dumped into a pipe hitting a target. The target is assumed to be a long thin graphite with dimension of 1.5 m x 8 mm (length x radius). The resulting hadrons were focused with six horn current settings ranging from 93 kA to 593 kA with a step of 100 kA. The two horn dimensions are assumed to be 4.2 m x 0.4 m and 3.8 m x 0.48 m (length x radius) respectively. After the horn focusing of the hadrons, the decay pipe is assumed to be $\sim$300 m long and 4 m in diameter to accommodate meson decays resulting in a neutrino beam having energy upto $\mathcal{O}(10\text{ GeV})$. Muon and hadron absorption is achieved using modular aluminum and steel absorber blocks, and the full system requires appropriate shielding and cooling. Two baselines are considered in this study, requiring adjustments to the horn orientation and neutrino beam angle to account for baseline length and Earth’s curvature. For reference, the LBNF beam is directed 5.8 below horizontal to reach the DUNE far detector at 1300 km. Using similar considerations, beam angles of approximately 4 and 13 below horizontal are obtained for the two proposed baselines in our study. A 1-ton WbLS near detector is proposed at a distance of 500 m to 1 km downstream of the decay pipe. Detailed optimization and engineering design are beyond the scope of this work.
@@
In Fig. 2, the simulated neutrino fluxes with 93 kA (red) and 493 kA (blue) are shown in the top left panel. The peaks match the first $\nu_{e}$ appearance peak locations. The top middle panel shows the neutrino interaction cross-section curves for various interaction channels. The top right panel shows the $\nu_{e}$ appearance oscillation probability for the two baselines. The bottom panels show the product of the flux and cross section, as well as the product of the flux, cross section and $\nu_{e}$ appearance oscillation probability. Each bottom panel corresponds to one interaction channel, namely charged-current inclusive, charged-current quasi-elastic and charged-current non-quasi-elastic from the left to right.
The CP violation phase sensitivity with our proposed detector depends on the solid curves. The final sensitivity includes the effect of absolute event rate, energy smearing and systematic uncertainties. It is obvious that for the 900 km baseline scenario, it is likely we can utilize only one appearance oscillation peak. On the contrary, for the 2900 km baseline scenario, we may use up to three oscillation peaks. The detector smearing may play an important to identify the oscillation peaks.
Figure 2: The top row shows the $\nu_{\mu}$ flux (simply abbreviated as F) $\Phi_{{\nu_{\mu}}}$; $\nu_{e}$ cross-section (X) $\frac{\sigma_{{\nu_{e}}}}{E}$; and the transition probabilities (P) $\text{P}_{\mu e}$ at EIC-SURF baseline of 2900 km, and at EIC-SNO baseline of 900 km. The bottom row shows the relevant products of these quantities (i.e., flux $\times$ cross-section or F.X, and flux $\times$ cross-section $\times$ probability or F.X.P) for three different types of interactions, -namely inclusive charge current ($C$), quasielastic charge current (CCQE), non-quasielastic charge current (CCNQE).